A company is projected to have a free cash flow of $243 million next year, growing at a 5.6% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.1%. The company's cost of capital is 12.7%. The company owes $71 million to lenders and has $17 million in cash. If it has 221 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.
Answer #1

A company is projected to have a free cash flow of $243 million next year, growing at a 5.6% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.1%. The company's cost of capital is 12.7%. The company owes $71 million to lenders and has $17 million in cash. If it has 221 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.

Projected Free Cash Flow (FCF1) = $ 243 million

Growth rate until year end 3 = 5.6%

growth rate thereafter (g1) = 2.1%

Cost of Capital (kc) = 12.7%

Calculating the Enterprise Value:

EV = EV = \frac{FCF_1}{(1+kc)^1}+\frac{FCF_1(1+g)}{(1+kc)^2}+\frac{FCF_1(1+g)^2}{(1+kc)^3}+\frac{FCF_1(1+g)^2(1+g_1)}{(1+kc)^3}*\frac{1}{(kc-g_1)}

EV = \frac{243}{(1+0.127)^1}+\frac{243(1+0.056)}{(1+0.127)^2}+\frac{243(1+0.056)^2}{(1+0.127)^3}+\frac{243(1+0.056)^2(1+0.021)}{(1+0.127)^3}*\frac{1}{(0.127-0.021)}

EV = $ 215.62 + $ 202.03 + $ 189.31 + $ 1823.40

EV = $ 2430.36 million

Enterprise Value = Value of Equity + Value of debt - Cash & Cash Equivalents

$ 2430.36 million = Value of Equity + $ 71 million - $ 17 million

Value of Equity = $ 2376.36 million

No of shares = 221 million

Value of Equity = Share price*No of shares

$ 2376.36 = Share Price*221 million shares

Share price = $ 10.8 per share